OpenStudy (anonymous):

On 1st Jan 2000, a man borrowed Rs. A at a fixed compound interest rate of a%. He pays back Rs. P on 1st Jan 2001, and then again Rs. P on 1st Jan 2002. Let a/100 = r Show that the amount still owing in Rs., just after this second payment, is A(1+r)^2 - P(1+r) - P.

hartnn (hartnn):

i am trying, after 1yr : in 2001 he needs to give A(1+r) pays back P so, A(1+r) -P now , you multiply [ A(1+r) -P] with 1+r for 2002 to get A(1+r)^2 -P(1+r) and again he pays back P, so final amount = A(1+r)^2 -P(1+r) -P phew...

OpenStudy (anonymous):

omg

OpenStudy (anonymous):

ok wait sub questions coming up bUT THANK YOU

OpenStudy (anonymous):

assuming tha the loan is completely paid back after N years, show that \[P = (Ar)/[1 - (1+r)^-N]\]

OpenStudy (anonymous):

ohh wait, i just substitute the sum formula, right?? for GP's??

hartnn (hartnn):

i guess yes, will have to try

OpenStudy (anonymous):

okie.

OpenStudy (anonymous):

ok, forget that other question (unless you have the answer, of course :P ) and just see this. Using the same P = blah bah walla formula, we, need to express N in terms of A, P, and r.

hartnn (hartnn):

means here ? \[P = (Ar)/[1 - (1+r)^{-N}]\] do u need to isolate N here ?

OpenStudy (anonymous):

yup.

hartnn (hartnn):

let me try (1-Ar)/P = (1+r)^(-N)

hartnn (hartnn):

logarithms use karna padega

OpenStudy (anonymous):

arrey no yaar. no other method??

OpenStudy (anonymous):

and btw no, it would be 1 - (Ar/P) = (1+r)^(-N)

hartnn (hartnn):

ha wohi :P

hartnn (hartnn):

without lod, N isolate nai ho sakta.

hartnn (hartnn):

*log

OpenStudy (anonymous):

well, ok.

hartnn (hartnn):

what u typing so much :P

hartnn (hartnn):

i think you can prove this by mathematical induction P=(Ar)/[1−(1+r)^(−N)] did u try it ?

OpenStudy (anonymous):

ummm......nope :P sorry, had gone for dinner.